Calculating the Direct Materials Mix Variance

Mangia Pizza Company makes frozen pizzas that are sold through grocery stores. Mangia developed the following standard mix for spreading on premade pizza shells to produce 16 giant-size sausage pizzas.

Direct Material Mix Mix Proportion SP Standard Cost
Tomato sauce 9.60 lbs 0.320 $1.50 $14.40
Cheese 10.80      0.360 2.70 29.16
Sausage 9.60      0.320 2.00 19.20
    Total 30.00 lbs $62.76

Mangia put a batch of 2,000 pounds of direct materials (enough for 1,067 frozen sausage pizzas) into process. Of the total, 700 pounds were tomato sauce, 900 pounds were cheese, and the remaining 400 pounds were sausage. The actual yield was 770 pizzas.

Required:

1. Calculate the standard mix (SM) in pounds for tomato sauce, for cheese, and for sausage.

SM
Tomato sauce ? pounds
Cheese ? pounds
Sausage ? pounds

2. Calculate the mix variance.
$ ? Unfavorable

3. Calculate the actual proportion used of tomato sauce, cheese, and sausage.

Actual mix proportion
Tomato sauce ? %
Cheese ? %
Sausage ? %
Answer #1

1. Calculation of standard mix in pounds:-

Total direct materials = 2000 pounds

Standard Ratio = .32 : .36 : .32

Revised Actual Quantity = Actual Quantity in Standard Ratio Proportion

Therefore, Standard Mix in pounds is as follows:-

Tomato Sauce : 2000 pounds * .32 = 640 pounds

Cheese : 2000 pounds * .36 = 720 pounds

Sausage : 2000 pounds * .32 = 640 pounds

2. Calculation of Mix Variance :-

Direct material Mix Variance = (Revised Actual Quantity - Actual Quantity) * Standard Price

RAQ AQ SP
Tomato Sauce 2000*.32=640 pounds 700 pounds $1.50
Cheese 2000*.36=720 pounds 900 pounds $2.70
Sausage 2000*.32=640 pounds 400 pounds $2.00
TOTAL 2000 pounds 2000 pounds

Tomato Sauce = (640-700)*$1.50 = $90 (Unfavourable)

Cheese = (720-900)*2.70 = $486 (Unfavourable)

Sausage = (640-400)*2 = $480 (Favourable)

Total = 90 (UNF) + 486 (UNF) + 480 (FAV) = 96 (Unfavourable)

3. Actual Proportion Used:

Tomato sauce = 700/2000*100 = 35 %

Cheese = 900/2000*100 = 45%

Sausage = 400/2000*100 = 20%